OpenStudy (anonymous):

]Which of the following is the slope between the points (0, -7) and (4, 12)? 19/4 -19/4 4/19 -4/19

Parth (parthkohli):

Okay, so slope is supposed to be change in y over change in x.

OpenStudy (anonymous):

yeah

Parth (parthkohli):

y goes 19 up and x goes 4 to the right

Parth (parthkohli):

Just make it into the form of a fraction and you'd get 19/4 :)

OpenStudy (anonymous):

thanksss. can u help me with this too Identify the correct slope and y intercept of the equation 3x - 3y = 4. slope = -1; y intercept at (0, -4/3) slope = 1; y intercept at (0, -4/3) slope = 3; y intercept at (0, -1) slope = -3; y intercept at (0, 1)

Parth (parthkohli):

* Slope - convert it into y = mx + b form and see what m is. * y intercept - convert it into y = mx + b form and see what's b.

OpenStudy (anonymous):

so its B right

Parth (parthkohli):

yes!

OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x